OpenStudy (imterribleatmath):

Given 2x+3y=18, find the equation of a line parallel to this line passing through (-3,-2)... I know that the slope is -2/3, in what formula do I plug in -3,-2 to get the answer/equation?

OpenStudy (anonymous):

Use this equation y-y1=m(x-x1) y1 is given, its -2 x1 is given, its -3 m=the slope is given as -2/3 Lets plug in and work it out y-(-2)=-2/3(x-(-3)) y+2=-2/3(x+3) distribute -2/3 y+2=-2/3x-2 -2 -2 _____________________________________________ y=-2/3x-4 ------> the equation